# FOF-X: Towards Real-time Detailed Human Reconstruction from a Single Image

### v3

@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on [1,1][-1,1] but not $[0,1]$, a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5).
Figure 4: The proposed Fourier Occupancy Field (FOF). Each pixel on the image II corresponds to a line in 3D space and can be described by an occupancy function $f(z)$, which can be expanded as Fourier series and represented with a vector $C(x,y)$. Therefore, the 3D human model can be encoded as a 2D vector field, forming our FOF. Note that the Fourier series here includes special forms such as the cosine series.
@@ III-B1 Reconstruction with Dual-sided Normal Maps
Thanks to the efficient dimensionality reduction with subspace approximation, the task of learning FOF from dual-sided normal maps can be essentially regarded as an image-to-image process. To this end, we exploit the image encoder built on the HR-Net [44] framework for its outstanding fitting capability in various vision tasks. We use the weak perspective camera model in our implementation, in which FOF (and thus the reconstructed geometry) is naturally aligned with the input RGB image. Note that FOF can also be used with the normalized device coordinate space (NDC space) so that various camera models, including the perspective camera, are compatible with the framework.
@@ III-B3 Supervision
To train our network, we sample points on the xyxy-plane and supervise the coefficients of the lines corresponding to these points with mean squared error (MSE) loss. To make the network more focused on human geometry, we only supervise the human foreground region of the image. The loss function $\mathcal{L}$ is formulated as:
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
To prepare training data, 3D meshes should be converted into FOF as training labels and input priors. By definition Eq. (5), we are to calculate the first few terms of the cosine series for each 1D occupancy signal $f(z)$. For generic signals, this would require numerical integration over a sampled version of $f(z)$, using e.g. the Adaptive Simpson Method, which however may introduce numerical errors and demand considerable computation.
Fortunately, thanks to the particular form of occupancy signals, the cosine coefficients can be derived analytically with exact solutions. Note that the 1D occupancy signal $f(z)$ is actually defined by discontinuities, where the line goes inside or outside the mesh. We extract the discontinuous points of $f(z)$ by designing a rasterizer-like procedure shown in Alg. 1. Formally, suppose that the line associated with $f(z)$ passes though the mesh for $k$ times. For the $i^{\textrm{th}}$ traversing, let $z_{i}$ and $z_{i}^{\prime}$ be the locations going inside and outside the mesh, respectively. Then, the set $\mathcal{Z}\triangleq\left\{(z_{i},z_{i}^{\prime})\right\}_{i=1}^{k}$ collects those intervals of the line that are inside the mesh. According to the definition of the occupancy signal, we have: $f(z)=1,z\in\mathcal{Z}$, and $f(z)=0$ otherwise, except for a finite number of discontinuities with measure zero. Therefore, we have the following analytical solution for the cosine coefficients:
@@
[…]
Algorithm 1 Mesh-to-FOF conversion
1:A triangle mesh $M=\{t_{i}\}_{i=1}^{n}$ $\triangleright$ A set of triangles
2:FOF $C\in\mathbb{R}^{H\times W\times N}$
[…]
6:for  $i=1$ to $n$ do $\triangleright$ parallelized
7:  Compute the bounding box $b_{i}$ of $t_{i}$
8:  Compute the normal direction $n_{i}$ of $t_{i}$
9:  for each pixel $(x,y)$ lying inside $b_{i}$ do
10:   if the pixel $(x,y)$ is inside $t_{i}$ then
11:     Compute the depth $z$ of $(x,y)$
12:     if $n_{i}$ faces front then
13:      $B[x,y]\leftarrow B[x,y]\cup{(z,0)}$ $\triangleright$ Add to buffer
14:     else
15:      $B[x,y]\leftarrow B[x,y]\cup{(z,1)}$ $\triangleright$ Add to buffer
16:     end if
17:   end if
18:  end for
19:end for
20:for $(x,y)\in\{1,...,H\}\times\{1,...,W\}$ do $\triangleright$ parallelized
21:  Sort the discontinuities in the buffer $B[x,y]$
22:  Apply the automaton to the buffer $B[x,y]$
23:  Compute integral results $\mathcal{I}$ with the buffer $B[x,y]$
24:  $C[x,y]\leftarrow\mathcal{I}$
25:end for
Figure 6: Example of Discontinuity Point Matching via Automaton.
However, due to numerical errors and some non-watertight meshes, rasterization may produce unpaired $\mathcal{Z}$, i.e., a line may encounter successive discontinuities both going out of or into the mesh. And it can come to a going outside point first instead of a going inside point. To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [19], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. Let’s consider the discontinuities corresponding to one pixel. We sort these points by $z$ first and then represent them with a sequence of 0s and 11s, where 0 stands for the normal vector facing front and 11 stands for the normal vector facing backwards. We add a stop letter \$\\$ to the sequence and then process it with the automaton shown in Fig. 5, which effectively handles irregular patterns (initial 11s and repeating 0s or 11s). After being filtered by the automaton, the produced sequence can then be integrated with Eq. (8). For watertight meshes, both our automaton and the conference version produce identical results, correctly converting meshes to FOF. For non-watertight meshes (e.g., surfaces with open boundaries or self-intersections, partly missing, or with non-manifold connectivity), the conference version method fails to compute FOF accurately, leading to floating artifacts (Fig. 2). Our automaton-based discontinuity matcher robustly handles irregular patterns arising from non-watertight meshes (including holes, self-intersections, and duplicate faces), effectively preventing reconstruction artifacts caused by such topological imperfections. Fig. 6 shows an example of our automaton processing a non-watertight mesh.
@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
Extracting meshes is necessary for applications such as mesh-based rendering and animation. Recent methods based on implicit neural representations need to apply an MLP network on a 3D sampling grid. Instead, as described in Eq. (4), the reconstruction of 3D occupancy field from FOF is simply a multiplication of two tensors. Then, the 3D mesh is extracted from the iso-surface of the 3D occupancy at the threshold of 0.5 with the Marching Cubes algorithm [45]. Both steps can be parallelized on GPUs for fast mesh generation.
@@
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and AA is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized. Unlike excessive Laplacian smoothing iterations that degrade geometric features into spherical approximations, our Laplacian coordinate constraint preserves topological fidelity and guarantees convergence to ground-truth mesh shapes.
@@
CAPE[46] THuman2.1[47] CustomHumans[48]
@@
PIFu[1] 2019 9.98 1.911 2.265 1.461 3.589 3.792 2.504 3.635 3.932 2.365
PIFuHD [11] 2020 15.90 3.279 3.393 1.903 3.963 3.986 2.620 4.184 4.115 2.432
PaMIR [12] 2021 36.48 1.426 1.629 1.377 1.549 1.771 1.720 1.211 1.404 1.350
ICON [3] 2022 35.00 0.967 1.001 0.909 0.968 0.881 1.338 0.798 0.755 1.039
ICONfl [3] 2022 35.56 0.817 0.814 0.755 0.948 0.860 1.254 0.811 0.761 0.984
D-IF [30] 2023 61.15 0.758 0.714 0.752 1.210 1.082 1.332 1.043 0.955 1.098
ECONif [4] 2023 124.55 0.947 0.956 0.928 1.186 1.224 1.442 1.015 1.027 1.106
ECONex[4] 2023 108.99 0.887 0.864 0.844 1.094 1.049 1.301 1.055 1.040 1.091
GTA [29] 2023 50.29 0.738 0.761 0.808 1.514 1.524 1.555 1.382 1.435 1.332
SiFU [2] 2024 62.38 0.676 0.677 0.733 1.015 1.004 1.271 0.880 0.900 0.994
SiTH [18] 2024 83.18 1.016 1.030 1.051 1.218 1.160 1.469 1.111 1.061 1.216
PSHuman [43] 2025 76.69 2.208 1.922 0.718 1.464 1.276 0.675 2.066 1.815 0.772
FOF-Base [19] 2022 0.09 4.583 3.825 2.345 3.597 3.408 2.489 4.677 4.084 2.671
FOF-SMPL [19] 2022 - 0.765 0.709 0.649 0.959 0.849 1.120 0.850 0.759 0.882
@@ IV-A Datasets
We collect THuman2.1 [47], CustomHumans [48], and CAPE [46] for our experiments. To maintain reproducibility, all the datasets we use are publicly accessible. For fair comparisons, we follow ICON [3] by training our method on THuman2.0, which is a subset of the first 526 items in THuman2.1, and using the CAPE dataset for evaluation. Additionally, we create two additional benchmarks on THuman2.1 and CustomHumans to evaluate our method and all the baselines.
@@
IV-B Implementation Details
The reconstruction network of FOF-X takes dual-sided normal maps and the FOF of SMPL as input. The dimensions of the input are $512\times 512\times(3+3+16)$, and the dimensions of the produced feature map are $128\times 128\times 256$. We adopt HRNet-W48 [49] as the image encoder of our system due to its multi-scale perception ability. Since the SMPL mesh only provides a rough shape prior, the high-frequency components corresponding to geometric details are redundant for the input. Empirically, we find that the first 16 terms of the SMPL mesh is sufficient for the shape prior. The decoder progressively upsamples the feature map to $256\times 256$ and $512\times 512$ using bilinear interpolations and residual convolutional blocks, while retaining a channel number of 256256 throughout the process. Finally, a convolutional layer projects the features to a FOF with a resolution of $512\times 512\times 128$. To cope with single view image as input, we use the off-the-shelf image-to-image translation network used in ICON [3] for generating dual-sided normal maps from the single image.
During training, generating multi-view normal maps and FOFs requires a considerable amount of storage and forwarding these data to GPUs can significantly slow down the training process. However, thanks to our fast FOF conversion algorithm implemented as a PyTorch CUDA extension, we can efficiently render meshes to FOF and normal maps. This allows us to keep all meshes in memory and generate training data on the GPU in real time. Additionally, this approach enables random view sampling during training, removing the constraints of training on fixed views.
We utilize the Adam optimizer [50] and set the learning rate to $2\times 10^{-5}$ without decay. Our methods and all the variants are trained on a single RTX-4090 GPU using automatic mixed-precision training with a batch size of 8. Convergence takes approximately 10 hours on the THuman2.0 dataset and about 2 days on the THuman2.1 training set.
For consistency with previous works [1, 4, 3], we employ a weak perspective camera. The scale of the weak perspective camera is fixed at 11 for training and evaluation, resulting in an orthogonal projection. But note that we can transform world space to NDC (Normalized Device Coordinates) space to support perspective cameras straightforwardly. We follow the post-processing for hands described in ECON[4], fusing the hands of the parametric model with the reconstructed body. This strategy is not used in our real-time systems.
IV-C Metrics
We use Chamfer distance, P2S (point-to-surface) distance, 3D IoU (Intersection over Union), and normal map error for quantitative evaluation. In the following, we provide a summary of metrics used in our experiments:
Chamfer and P2S Distances
We use Chamfer and P2S distances to evaluate the surface geometric error between the predicted and ground-truth meshes. Here, the Chamfer distance $d_{cd}$ is defined as:
$d_{cd}(M_{pr},S_{pr},M_{gt},S_{gt})=\frac{1}{2\left|S_{pr}\right|}$ $\sum_{x\in P_{pr}}d(x,M_{gt})+$ (10)
$\frac{1}{2\left|S_{gt}\right|}$ $\sum_{x\in S_{gt}}d(x,M_{pr}),$
where $S_{pr},S_{gt}$ are the sets of points uniformly sampled on the predicted and ground-truth mesh surfaces $M_{pr}$ and $M_{gt}$ separately, and $d(x,M)$ is the minimal distance from a point $x$ to the mesh surface $M$. We report the averaged Chamfer distances of all test samples. We also measure the average point-to-surface distance from the points on the reconstructed surface to the ground-truth mesh. The P2S distance $d_{p2s}$ is defined as:
$d_{p2s}(S_{pr},M_{gt})=\frac{1}{\left|S_{pr}\right|}\sum_{x\in P_{pr}}d(x,M_{gt}).$ (11)
Note that this metric alone does not fully measure geometric accuracy. For instance, the P2S distance between a ground-truth mesh and a partial ground-truth mesh would be zero. However, we also report this metric to align with the existing published works.
Normal Difference
To measure the visual quality of the reconstructed geometries, we render them as normal maps and compute the mean squared error with the ground-truth normal maps. For each reconstructed mesh, we render four normal maps from views at $\{0^{\circ},90^{\circ},180^{\circ},270^{\circ}\}$ based on vertex normals. We do not apply any anti-aliasing techniques when rendering these normal maps. Such a metric provides an effective assessment of the visual fidelity of the reconstructions.
@@ IV-C Metrics
IV-D Comparison with the State of the Arts
We compare our method with the state-of-the-art approaches, including PIFu [1], PIFuHD [11], PaMIR[12], ICON [3], ECON[4], D-IF[30], GTA[29], SiFU[2], SiTH[18], and PSHuman[43] on CAPE[46], THuman2.1[47], and CustomHumans[48].
@@ IV-D Comparison with the State of the Arts
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[3]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[11] and PSHuman[43], are trained on THuman2.0[47]. It is worth noting that PSHuman’s[43] dataset includes THuman 2.1. For PIFu[1] and PaMIR[12], similar to ICON[3], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $512\times 512\times 512$ resolution for the input of all these methods. We do not compare our method with [13] and [51] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
As can be seen, our methods show the lowest reconstruction errors compared to other methods on CustomHumans datasets. As mentioned in [18], CAPE provides the fitted SMPL+D meshes instead of the original scanned meshes as the ground-truth, leading to mismatches between the image details and the ground-truth meshes. This may cause a potential bias in evaluation. Despite this, our FOF-X can still achieve a comparable metric on CAPE. In addition, on the THuman2.1 dataset, we are only slightly inferior to PSHuman[43] in terms of normals. However, we note that PSHuman’s[43] training dataset incorporates THuman2.1 without providing clear train/test splits, potentially allowing some test data to be included in training. Notably, our method demonstrates robustness across all benchmarks, achieving better or comparable performance to other computationally heavy methods.
@@
We conduct a qualitative comparison in Fig. 8. As can be seen, SiTH[18] and PSHuman[43] generally produces good results but sometimes introduces artifacts in the limbs and side shapes. ECON[4] suffers from self-occlusions which is the essential limitation of its optimization-based solution; SiFU[2] cannot recover detailed surfaces, although it can produce accurate rough shapes; D-IF[30] has difficulties dealing with limb shapes due to the way it uses the SMPL prior. In contrast, our method is able to reconstruct plausible 3D human models under different body poses and shapes. In terms of surface quality and geometric accuracy, our method outperforms other state-of-the-art approaches. Notably, our method is much more lightweight than others, allowing for real-time reconstruction.
Fig. 9 shows additional comparison on non-frontal view inputs. Generally, it is more challenging to obtain information about the body and limb shape from non-frontal view inputs. In this case, the results of SiTH[18], D-IF[30], and ECON[4] all experience varying degrees of performance degradation. SiFU[2] is still constrained by the resolution issue in producing detailed surfaces. PSHuman[43] has certain difficulties in processing limb shapes. In contrast, our method is more robust and still achieves good results.
Figure 9: Qualitative comparison against the state-of-the-art methods on non-frontal view inputs.
@@
Table II shows the comparison results in terms of running time. All methods except PSHuman[43] are tested on a single RTX-3090 GPU, where FOF-X represents the acceleration of FOF-X using TensorRT. PSHuman[43] is tested on the A100 GPU due to memory limitations. Our FOF-X takes 0.02 seconds per frame, demonstrating 500× acceleration over the fastest baseline (PIFu[1]: 9.98s). The original system presented in our conference publication, FOF[19], is implemented in Python. Due to limitations in the original mesh-to-FOF algorithm, the SMPL prior is not supported in the FOF real-time system. Since FOF-SMPL requires the execution of three separate scripts, its running time is not reported. Our enhanced pipeline FOF-X adds additional components such as SMPL estimation, image segmentation, and rendering while maintaining real-time performance. Notably, all variants of our method (Python/TensorRT) consistently outperform existing works in terms of efficiency by two orders of magnitude.
IV-E Ablation Study

### v4

@@ III-A2 FOF with Cosine Series
Because $f(z)$ is defined on [1,1][-1,1] but not $[0,1]$, a shift on $z$ is needed, which contributes to the term $\frac{(z+1)}{2}$ in Eq. (5). We further analyze the representation efficiency of Fourier and cosine series under the same FOF formulation. As shown in the supplementary material, cosine series consistently achieve lower approximation error with fewer coefficients, motivating our design choice.
@@ III-B1 Reconstruction with Dual-sided Normal Maps
Thanks to the efficient dimensionality reduction with subspace approximation, the task of learning FOF from dual-sided normal maps can be essentially regarded as an image-to-image process. To this end, we exploit the image encoder built on the HR-Net [WangSCJDZLMTWLX19] framework for its outstanding fitting capability in various vision tasks. We use the weak perspective camera model in our implementation, in which FOF (and thus the reconstructed geometry) is naturally aligned with the input RGB image. Note that FOF can also be used with the normalized device coordinate space (NDC space) so that various camera models, including the perspective camera, are compatible with the framework.
@@ III-B3 Supervision
To train our network, we sample points on the $xy$-plane and supervise the coefficients of the lines corresponding to these points with mean squared error (MSE) loss. To make the network more focused on human geometry, we only supervise the human foreground region of the image. The loss function $\mathcal{L}$ is formulated as:
@@
Figure 5: Example of Discontinuity Point Matching via Automaton.
@@ III-C1 Mesh to FOF with Automaton-based Discontinuity Matcher
To prepare training data, 3D meshes are converted into FOF as training labels and input priors. According to Eq. (5), this amounts to computing the first few terms of the cosine series for each 1D occupancy signal $f(z)$. Unlike generic signals requiring numerical integration, occupancy signals allow exact analytic solutions.
Formally, suppose the line associated with $f(z)$ intersects the mesh $k$ times, with $(z_{i},z_{i}^{\prime})$ denoting the entry and exit points. Then $\mathcal{Z}={(z_{i},z_{i}^{\prime})}{i=1}^{k}$ defines the inside intervals, and the cosine coefficients are
@@
where $t_{n}=\tfrac{n\pi}{2}$. Since $k$ is typically small (1–3 for human meshes), each coefficient can be computed in $O(1)$, and the whole mesh-to-FOF conversion runs in $O(HWN)$ with GPU parallelization, where $H,W$ are FOF resolution.
In practice, ground-truth meshes are not always watertight, leading to irregular discontinuity sequences (e.g., consecutive entries or exits). To solve this, we propose an automaton-based discontinuity matcher to enhance the robustness. Unlike what we do in the conference version [fof], we collect those discontinuities and normal vectors for each line first and process them with an automaton before working out the integral. For watertight meshes, both our automaton and the conference version produce identical results, correctly converting meshes to FOF. For non-watertight meshes (e.g., surfaces with open boundaries or self-intersections, partly missing, or with non-manifold connectivity), the conference version method fails to compute FOF accurately, leading to floating artifacts (Fig. 2). Our automaton-based discontinuity matcher robustly handles irregular patterns arising from non-watertight meshes (including holes, self-intersections, and duplicate faces), effectively preventing reconstruction artifacts caused by such topological imperfections. We use the automaton described in Fig. 4 to handle all the lights. For a more detailed approach please refer to section A of the supplementary material. Fig. 5 shows an example of our automaton processing a non-watertight mesh.
@@ III-C2 FOF to Mesh with Laplacian Coordinate Constraint
Extracting meshes is necessary for applications such as mesh-based rendering and animation. Recent methods based on implicit neural representations need to apply an MLP network on a 3D sampling grid. Instead, as described in Eq. (4), the reconstruction of 3D occupancy field from FOF is simply a multiplication of two tensors. Then, the 3D mesh is extracted from the iso-surface of the 3D occupancy at the threshold of 0.5 with the Marching Cubes algorithm [marching]. Both steps can be parallelized on GPUs for fast mesh generation.
@@
where $D$ is a diagonal matrix with $D_{ii}=d_{i}$ on the diagonal, $d_{i}$ is the number of neighbors of the $i^{\textrm{th}}$ point, $X\in\mathbb{R}^{n\times 3}$ is the coordinate matrix to be optimized, and $A$ is the adjacency matrix. $A_{ij}=1$ if points $v_{i}$ and $v_{j}$ are adjacent, otherwise $A_{ij}=0$. $U$ is the index set of unreliable points, which are to be optimized. Unlike excessive Laplacian smoothing iterations that degrade geometric features into spherical approximations, our Laplacian coordinate constraint preserves topological fidelity and guarantees convergence to ground-truth mesh shapes.
@@
CAPE[cape1] THuman2.1[tao2021function4d] CustomHumans[cus]
@@
PIFu[saito2019pifu] 2019 9.98 1.911 2.265 1.461 3.589 3.792 2.504 3.635 3.932 2.365
PIFuHD [saito2020pifuhd] 2020 15.90 3.279 3.393 1.903 3.963 3.986 2.620 4.184 4.115 2.432
PaMIR [zheng2021pamir] 2021 36.48 1.426 1.629 1.377 1.549 1.771 1.720 1.211 1.404 1.350
ICON [xiu2022icon] 2022 35.00 0.967 1.001 0.909 0.968 0.881 1.338 0.798 0.755 1.039
ICONfl [xiu2022icon] 2022 35.56 0.817 0.814 0.755 0.948 0.860 1.254 0.811 0.761 0.984
D-IF [yang2023dif] 2023 61.15 0.758 0.714 0.752 1.210 1.082 1.332 1.043 0.955 1.098
ECONif [econ] 2023 124.55 0.947 0.956 0.928 1.186 1.224 1.442 1.015 1.027 1.106
ECONex[econ] 2023 108.99 0.887 0.864 0.844 1.094 1.049 1.301 1.055 1.040 1.091
GTA [gta] 2023 50.29 0.738 0.761 0.808 1.514 1.524 1.555 1.382 1.435 1.332
SiFU [sifu] 2024 62.38 0.676 0.677 0.733 1.015 1.004 1.271 0.880 0.900 0.994
SiTH [sith] 2024 83.18 1.016 1.030 1.051 1.218 1.160 1.469 1.111 1.061 1.216
PSHuman [li2025pshuman] 2025 76.69 2.208 1.922 0.718 1.464 1.276 0.675 2.066 1.815 0.772
FOF-Base [fof] 2022 0.09 4.583 3.825 2.345 3.597 3.408 2.489 4.677 4.084 2.671
FOF-SMPL [fof] 2022 - 0.765 0.709 0.649 0.959 0.849 1.120 0.850 0.759 0.882
@@ IV-A Datasets
We collect THuman2.1 [tao2021function4d], CustomHumans [cus], and CAPE [cape1] for our experiments. To maintain reproducibility, all the datasets we use are publicly accessible. For fair comparisons, we follow ICON [xiu2022icon] by training our method on THuman2.0, which is a subset of the first 526 items in THuman2.1, and using the CAPE dataset for evaluation. Additionally, we create two additional benchmarks on THuman2.1 and CustomHumans to evaluate our method and all the baselines.
@@
IV-B Comparison with the State of the Arts
We compare our method with the state-of-the-art approaches, including PIFu [saito2019pifu], PIFuHD [saito2020pifuhd], PaMIR[zheng2021pamir], ICON [xiu2022icon], ECON[econ], D-IF[yang2023dif], GTA[gta], SiFU[sifu], SiTH[sith], and PSHuman[li2025pshuman] on CAPE[cape1], THuman2.1[tao2021function4d], and CustomHumans[cus].
@@ IV-B Comparison with the State of the Arts
The quantitative results are summarized in Table II. denotes the re-implemented version in ICON[xiu2022icon]. denotes the re-implemented version with the cosine series and the improved inter-conversion algorithms. FOF-X(HR32) and FOF-X(HR48) are FOF-X model with HRNet-W32 and HRNet-W48 respectively. All methods listed, except PIFuHD[saito2020pifuhd] and PSHuman[li2025pshuman], are trained on THuman2.0[tao2021function4d]. It is worth noting that PSHuman’s[li2025pshuman] dataset includes THuman 2.1. For PIFu[saito2019pifu] and PaMIR[zheng2021pamir], similar to ICON[xiu2022icon], we use the re-implemented versions, as the official releases are trained with additional commercial datasets. Since some methods do not provide the training sample lists, we use THuman2.1 instead of THuman2.0 to evaluate all methods. We use $512\times 512\times 512$ resolution for the input of all these methods. We do not compare our method with [alldieck2022phorhum] and [He2021ARCHAC] because their codes are not released. For simplicity, we also ignore some works that have already been compared with the state-of-the-art methods listed above. We found that some methods are sensitive to the ratio of the human subject’s size in the image. To ensure consistency, we use a ratio of 0.9 for all methods, corresponding to the 1.8m height normalization mentioned in the datasets section. Ground-truth SMPL meshes are used to better demonstrate the capability of the evaluated methods.
As can be seen, our methods show the lowest reconstruction errors compared to other methods on CustomHumans datasets. As mentioned in [sith], CAPE provides the fitted SMPL+D meshes instead of the original scanned meshes as the ground-truth, leading to mismatches between the image details and the ground-truth meshes. This may cause a potential bias in evaluation. Despite this, our FOF-X can still achieve a comparable metric on CAPE. In addition, on the THuman2.1 dataset, we are only slightly inferior to PSHuman[li2025pshuman] in terms of normals. However, we note that PSHuman’s[li2025pshuman] training dataset incorporates THuman2.1 without providing clear train/test splits, potentially allowing some test data to be included in training. Notably, our method demonstrates robustness across all benchmarks, achieving better or comparable performance to other computationally heavy methods.
@@
We conduct a qualitative comparison in Fig. 7. As can be seen, SiTH[sith] and PSHuman[li2025pshuman] generally produces good results but sometimes introduces artifacts in the limbs and side shapes. ECON[econ] suffers from self-occlusions which is the essential limitation of its optimization-based solution; SiFU[sifu] cannot recover detailed surfaces, although it can produce accurate rough shapes; D-IF[yang2023dif] has difficulties dealing with limb shapes due to the way it uses the SMPL prior. In contrast, our method is able to reconstruct plausible 3D human models under different body poses and shapes. In terms of surface quality and geometric accuracy, our method outperforms other state-of-the-art approaches. Notably, our method is much more lightweight than others, allowing for real-time reconstruction.
@@
Table II shows the comparison results in terms of running time. All methods except PSHuman[li2025pshuman] are tested on a single RTX-3090 GPU, where FOF-X represents the acceleration of FOF-X using TensorRT. PSHuman[li2025pshuman] is tested on the A100 GPU due to memory limitations. Our FOF-X takes 0.02 seconds per frame, demonstrating 500× acceleration over the fastest baseline (PIFu[saito2019pifu]: 9.98s). The original system presented in our conference publication, FOF[fof], is implemented in Python. Due to limitations in the original mesh-to-FOF algorithm, the SMPL prior is not supported in the FOF real-time system. Since FOF-SMPL requires the execution of three separate scripts, its running time is not reported. Our enhanced pipeline FOF-X adds additional components such as SMPL estimation, image segmentation, and rendering while maintaining real-time performance. Notably, all variants of our method (Python/TensorRT) consistently outperform existing works in terms of efficiency by two orders of magnitude.
IV-C Ablation Study